OpenStudy (firejay5):

12. Suppose you drive in a straight line 10 km north from home, then turn west and drive 12 km in a straight line. How far are you from home?

OpenStudy (firejay5):

@anonymous_user

OpenStudy (anonymous_user):

22 km, I believe.

OpenStudy (firejay5):

I got 15.6km, but I am not for sure. How did you get 22km

OpenStudy (anonymous_user):

12 + 10 = 22

OpenStudy (firejay5):

I thought it would Pythagorean theorem

OpenStudy (anonymous_user):

I'm not sure.

OpenStudy (firejay5):



OpenStudy (nincompoop):

are you even trying to solve the questions you're posting?

OpenStudy (firejay5):

quit accusing me dude/girl geez louise

OpenStudy (nincompoop):

are you?

OpenStudy (firejay5):

yes DO YOU SEE MY posts

OpenStudy (nincompoop):

where is your solution then? are you just waiting for an answer or clue?

OpenStudy (firejay5):

I need a answer to support my reasoning

OpenStudy (nincompoop):

where is your reasoning? what are the steps have you taken?

OpenStudy (nincompoop):

here's your pythagorean theorem \( \sf r^2 = x^2 + y^2 \)

OpenStudy (amistre64):

your directions are relative, but the reasoning is sound. pythag or vectors to determine displacement

OpenStudy (nincompoop):

here's a quick glance when solving for resolution of vectors (resultant r)